# Permutation-based multiple testing when fitting many generalized linear models

### v1

@@ 3 Multiple responses
Let $\bm{M}^{n}_{0}$ be the $w$-by-$m$ matrix with $(j,l)$-th entry equal to $S^{*,j,l}_{\bm{\gamma}_{0}}$. Note that $\bm{M}^{n}_{0}$ is based on knowledge of the true nuisance parameters $\bm{\gamma}_{0}$. The consequence is that each entry of the matrix $\bm{M}^{n}_{0}$ is the sum of $n$ independent (flipped) score contributions. Further, note that each row of $\bm{M}^{n}_{0}$ is uncorrelated with the other rows, due to the independence of the flips. Finally, the correlation structure within each row coincides with the correlation structure of the contributions $\nu^{*1}_{\bm{\gamma}_{0}},\dots,\nu^{*m}_{\bm{\gamma}_{0}}$. Consequently, the multivariate central limit theorem (Van der Vaart, 1998) implies that $\bm{M}^{n}_{0}$ converges in distribution to some matrix $\bm{M}_{0}$, which has identically distributed multivariate normal rows.
Now it is left to show that $\bm{M}$, i.e., the matrix based on estimated nuisance parameters, also has identically distributed multivariate normal rows. As shown in the proof of Theorem 2 in Hemerik et al. (2020), we have $S^{*,j,l}_{\hat{\bm{\gamma}}}=S^{*,j,l}_{\bm{\gamma}_{0}}+o_{p}(1)$, $1\leq j\leq w$, $1\leq l\leq m$. This means that $\bm{M}$ is asymptotically equivalent to $\bm{M}_{0}$. Hence the result holds. ∎
If instead of effective scores we used standardized effective scores to fill the matrix $\bm{M}^{n}$, then Lemma 4 will still hold, since the standardized effective scores are asymptotically equivalent to the unstandardized effective scores. This is detailed in De Santis et al. (2022).
A special case of interest relates to homoscedastic linear regression models. The matrix of weights becomes an identity matrix, which causes some simplifications in the formulas; in particular
$S^{*,j,l}_{\hat{\bm{\gamma}}}=X^{\prime}(I-{H})G_{j}(Y^{l}-\hat{\mu}^{l})$
and
$Var\left\{S^{*,j,l}_{\hat{\bm{\gamma}}}\right\}=X^{\prime}(I-{H})G_{j}(I-{H})G_{j}(I-{H})X.$
while the standardized score is still
$S^{S,j,l}_{\hat{\bm{\gamma}}}=S^{*,j,l}_{\hat{\bm{\gamma}}}Var\left\{S^{*,j,l}_{\hat{\bm{\gamma}}}\right\}^{-1/2}.$
For the linear regression model, using the standardized scores, we are able to get finite sample results, as stated in the following Lemma. This reflects that we still get the second-moment null-invariance property for linear regression models with normal response, while it is not true in general.
Lemma 5.
Assume a multivariate linear regression model with normal response, where the covariance between responses is equal among different units, that is,
$E[(Y^{l}-\mu^{l})(Y^{p}-\mu^{p})^{\prime}]=\sigma_{pl}\bm{I}_{n}$
where $\bm{I}_{n}$ is the $n$-dimensional identity matrix. Let $\bm{M}^{S,n}$ be the $w$-by-$m$ matrix with $(j,l)$-entry equal to $S^{S,j,l}_{\hat{\bm{\gamma}}}$. Then $\bm{M}^{S,n}$ has independent rows with the same multivariate normal distribution.
@@
The rows of the matrix $\bm{M}^{S,n}$ are uncorrelated, due to the independence of the flips. Within the row the correlation structure is
$Cov\left\{S^{S,j,l}_{\hat{\bm{\gamma}}},S^{S,j,p}_{\hat{\bm{\gamma}}}\right\}=Var\left\{S^{*,j,l}_{\hat{\bm{\gamma}}}\right\}^{-1/2}E\left[S^{*j,l}_{\hat{\bm{\gamma}}}\left(S^{*,j,p}_{\hat{\bm{\gamma}}}\right)^{\prime}\right]Var\left\{S^{*,j,p}_{\hat{\bm{\gamma}}}\right\}^{-1/2}$
Using the result
$E[(Y^{l}-\mu^{l})(Y^{p}-\mu^{p})^{\prime}]=\sigma_{pl}\bm{I}_{n}$
we get, after simple computations,
$Cov\left\{S^{S,j,l}_{\hat{\bm{\gamma}}},S^{S,j,p}_{\hat{\bm{\gamma}}}\right\}=\sigma_{pl}/(\sigma_{p}\sigma_{l}),$
which is independent of the flip, where $\sigma_{l}$ is the variance of the $l$-th response. It follows that the matrix $M^{S,n}$ has independent and identically distributed multivariate normal rows for finite sample size. ∎
From Lemma 4 (or 5) we can build an asymptotically exact local $\alpha$-level test for any composite hypothesis as follows. Here we derive the results for a two-sided alternative. Let $H_{0}$ as defined in (3). Define $\psi:\mathbb{R}^{k}\rightarrow\mathbb{R}$ as a function non decreasing in its argument. We can derive a global flipped test statistic as
$T^{n}_{j}=\psi\left(|S^{S,j,1}_{\hat{\gamma}}|,\dots,|S^{S,j,L}_{\hat{\gamma}}|\right).$ (5)
The following theorem shows that from this test statistic we can get an asymptotic $\alpha$-level test.
@@
For every $1\leq j\leq w$, consider the statistic $T_{j}^{n}$ as defined in (5) and let $T^{n}_{(1)}\leq...\leq T^{n}_{(w)}$ be the sorted statistics. Consider the test that rejects if $T_{1}^{n}>T^{n}_{\lceil(1-\alpha)w\rceil}$. As $n\rightarrow\infty$, under $H_{0}$ the rejection probability converges to $\lfloor\alpha w\rfloor/w\leq\alpha$.
@@
Lemma 4 implies that the test statistics $T^{n}_{1},\dots,T^{n}_{w}$ are asymptotically independent and identically distributed. Note that, by the definition of $\psi$, high values of $T_{1}$ shows evidence against the null hypothesis $H$. Hence, Lemma 1 of Hemerik et al. (2020) directly applies and we derive that we get an asymptotically $\alpha$-level test. ∎
Theorem 6 implies that we can build an asymptotic valid local test for the hypothesis (3), which therefore guarantees weak control of the FWER. Note that for the linear model we can obtain finite sample results, because we can directly apply Lemma 5 and it follows that the test has finite sample size properties.
Theorem 7.
Assume to fit a multivariate linear regression model with normal response. For every $1\leq j\leq w$, consider the statistic $T_{j}^{n}$ as defined in (5) and let $T^{n}_{(1)}\leq...\leq T^{n}_{(w)}$ be the sorted statistics. Consider the test that rejects if $T_{1}^{n}>T^{n}_{\lceil(1-\alpha)w\rceil}$. Under $H_{0}$ the rejection probability is $\leq\alpha$.
Proof.
Lemma 5 implies that the test statistics $T^{n}_{1},\dots,T^{n}_{w}$ are independent and identically distributed. Note that, by the definition of $\psi$, high values of $T_{1}$ shows evidence against the null hypothesis $H$. Hence, Theorem 1 of Hemerik and Goeman (2018) directly applies and the test derived is an $\alpha$-level test for finite sample size. ∎
Multiple choices of the function $\psi$ are available (Pesarin, 2001) and the choice will influence the power properties in different settings. We can subsequently apply the closed testing approach (Marcus et al., 1976) to build a procedure which guarantees strong control of the FWER by computing the $2^{n}$ intersection tests, and this procedure is optimal in the sense that every multiple testing procedure is equal or can be improved by applying the closed testing principle (Goeman et al., 2021) .
However, the number of tests might be unfeasible for large values of $m$. A dramatic shortcut is given by selecting the maximum of the test statistics as combining function in the following way. This is called the max-$T$ approach. For sake of completeness, we will give a direct proof of its validity.
There are roughly two versions of the max-$T$ method by Westfall and Young (Westfall and Young, 1993; Westfall and Troendle, 2008; Meinshausen et al., 2011): the single-step method and the sequential method. The single-step method is simpler and faster, while the sequential method is more powerful. The single-step max-$T$ method, based on the matrix $\bm{M}^{n}$ of test statistics, is defined as follows. Here we formulate a version that employs two-sided tests. For every $1\leq j\leq w$, let $m_{j}$ be the maximum of the test statistics $|\bm{M}^{n}_{j,l}|$, $1\leq l\leq m$. Let $m_{(1)},\dots,m_{(w)}$ be the sorted values $m_{1},\dots,m_{w}$. Then the multiple testing method rejects all hypotheses with index $l$ for which $|\bm{M}^{n}|_{1,l}>m_{(\lceil(1-\alpha)w\rceil)}$. The sequential max-$T$ method is defined as follows; after the first step defined above is completed, the procedure is continued in a step-down way. We remove from the matrix $\bm{M}^{n}$ all rows corresponding to the hypotheses rejected in the first step, then we apply again the same procedure described above. The procedure can be continued until we have no more rejections. Note that the test with standardized test statistic is defined in the same way.
The following theorem states that the single-step and sequential max-$T$ methods provide strong asymptotic FWER control. Write $FWER_{n}$ to indicate potential dependence of the FWER on $n$.
Theorem 8.
For both the single-step and sequential max-$T$ method, $\limsup_{n\rightarrow\infty}(FWER_{n})\leq\alpha.$
Proof.
For both the single-step and the sequential max-$T$ method, the argument is as follows. Recall that $\bm{M}^{n}$ converges in distribution to $\bm{M}$. Let $\bm{M}_{\mathcal{N}}$ be the submatrix of $\bm{M}$ that only contains the rows corresponding to the true hypotheses. If the matrix $\bm{M}$ is used as input for the multiple testing procedure, then strong FWER control follows directly from the fact that the rows of $\bm{M}_{\mathcal{N}}$ are exchangeable, i.e. swapping rows does not change the distribution of the matrix.
If instead $\bm{M}^{n}$ is used as input, then FWER control follows from the continuous mapping theorem, since $\bm{M}^{n}$ converges to $\bm{M}$ in distribution. This finishes the proof.
Finally, note that for the procedure that uses standardized scores for linear regression model the control of the familywise error rate is obtained for finite sample size, as the matrix $\bm{M}^{S,n}$ has an exact multivariate normal distribution. ∎
@@ 4.1 Univariate
We first show a simulation study for univariate tests in Figure 1. We test $H_{0}:\beta=0$ setting the regression parameters $(\beta,\gamma)=(0,1).$ The correlation between covariates is equal to $0.5$ while a total of 100 000 simulations have been run. We compare the Flipscores approach with the three standard parametric competitors, namely the Wald, Score and Likelihood Ratio (LRT) tests. The Flipscores test is based on 2 000 random flips (we refer to Hemerik and Goeman (2018) for a discussion about the use of a limited number of random flips). The $x$ axis represents the number of hypotheses tested, while the $y$ axis represents the ratio between the Empirical type I error and the nominal level $\alpha$. A Bonferroni correction is applied to the number of hypotheses tested.
We can observe that the Flipscores is always inside the 95% simulation confidence bands. The Score test is slightly conservative while the two other parametric methods are less satisfactory for opposite reasons. More simulations can be find in the appendix for different sample size, value of nuisance coefficient and correlation.
Figure 1: Logit model, univariate. On abscissa the Bonferroni-adjusted alpha for the control of FWER at level 5%, i.e. .05/m
@@ 4.2 Multivariate
Figures 2 and 3 represent a multivariate simulation. We set a total of 1 000 dependent variables. For 20% they have $\beta=1$ and remaining 80 % have $\beta=0$. We set $\gamma=-1$, and the correlation between covariates equal to 0.5. Again, 2 000 random flips are used.
For controlling the FWER the proposed method is used for the Flipscores approach, while the other parametric competitors are corrected with the Bonferroni-Holm procedure.
The $x$ axis of both figures represents the average observed correlation between the dependent variables. Figure 2 represents the empirical FWER of the true null hypothesis. The flipscores is closed to the nominal level, the LRT is anticonservative for low correlation, while the other two methods are highly conservative. Figure 3 represents the empirical power for the true alternatives. The flipscores has greater power, especially for higher correlation, while the other methods do not take advantage of the correlation structure. We observe similar results for different settings in the appendix. Indeed, the proposed method seems to be more satisfactory especially when high correlation is present.
Figure 2: Logit model, multivariate
Figure 3: Logit model, multivariate

### v2

@@ 3 Multiple dependent variables
Let $\bm{M}^{n}_{0}$ be the $w$-by-$m$ matrix with $(j,l)$-th entry equal to $S^{*,j,l}_{\bm{\gamma}_{0}}$. Note that $\bm{M}^{n}_{0}$ is based on knowledge of the true nuisance parameters $\bm{\gamma}_{0}$, where $S^{*,j,l}_{\bm{\gamma}_{0}}$ is the corresponding effective score. The consequence is that each entry of the matrix $\bm{M}^{n}_{0}$ is the sum of $n$ independent (flipped) score contributions (Hemerik et al., 2020). In particular each row of $\bm{M}^{n}_{0}$ is uncorrelated with the other rows, due to the independence of the flips. Moreover, the correlation structure within each row coincides with the correlation structure of the contributions $\nu^{*1}_{\bm{\gamma}_{0}},\dots,\nu^{*m}_{\bm{\gamma}_{0}}$. This follows as, the sum of $n$ independent vectors with the same correlation structure has the same correlation structure as the summands, and dividing by $n^{1/2}$ leaves the correlation structure unchanged. Consequently, the multivariate central limit theorem (Van der Vaart, 1998) implies that $\bm{M}^{n}_{0}$ converges in distribution to some matrix $\bm{M}_{0}$, which has identically distributed multivariate normal rows.
Now it remains to show that $\bm{M}$, i.e., the matrix based on the estimated nuisance parameters, also has identically distributed multivariate normal rows. As shown in the proof of Theorem 2 in Hemerik et al. (2020), we have $S^{*,j,l}_{\hat{\bm{\gamma}}}=S^{*,j,l}_{\bm{\gamma}_{0}}+o_{p}(1)$, $1\leq j\leq w$, $1\leq l\leq m$. This means that $\bm{M}$ is asymptotically equivalent to $\bm{M}_{0}$. Hence the result holds. ∎
If we used standardized scores, instead of effective scores, to fill the matrix $\bm{M}^{n}$, then Lemma 3 still holds, since the standardized scores are asymptotically equivalent to the effective (unstandardized) scores, see (De Santis et al., 2022).
From Lemma 3 we can build an asymptotic exact $\alpha$-level test for any composite hypothesis as follows. Let $H_{L}$ be defined as in (3). Given any non-decreasing function $\psi:\mathbb{R}^{k}\rightarrow\mathbb{R}$, for each $1\leq j\leq w$, define global flipped test-statistics
$T^{n}_{j}=\psi\left(|S^{j,1}_{\hat{\gamma}}|,\dots,|S^{j,L}_{\hat{\gamma}}|\right).$ (5)
The following theorem shows that we can use these test statistics in order to obtain an asymptotic $\alpha$-level test.
Theorem 4.
For every $1\leq j\leq w$, consider the statistic $T_{j}^{n}$ as defined in (5) and let $T^{n}_{(1)}\leq...\leq T^{n}_{(w)}$ be the sorted statistics. Consider the test that rejects if $T_{1}^{n}>T^{n}_{\lceil(1-\alpha)w\rceil}$. As $n\rightarrow\infty$, under $H_{L}$ the rejection probability converges to $\lfloor\alpha w\rfloor/w\leq\alpha$.
@@
Lemma 3 implies that the test statistics $T^{n}_{1},\dots,T^{n}_{w}$ are asymptotically independent and identically distributed. Note that, by the definition of $\psi$, high values of $T_{1}$ shows evidence against the null hypothesis $H_{L}$. Hence, Lemma 1 of Hemerik et al. (2020) directly applies and obtain anasymptotic $\alpha$-level test. ∎
Theorem 4 implies that we can build an asymptotic valid local test for the hypothesis (3), which therefore guarantees weak control of the familywise error rate (FWER), that is, the probability of making any false rejection under the global null hypothesis.
Multiple choices of the function $\psi$ are available (Pesarin, 2001) and the choice will influence the power properties in different settings. We can subsequently apply the closed testing approach (Marcus et al., 1976) to build a procedure which guarantees strong control of the FWER by computing the $2^{n}$ intersection tests, and this procedure is optimal in the sense that every FWER controlling procedure is equal or can be improved by applying the closed testing principle (Goeman et al., 2021).The strong control of the FWER ensures control of the probability of making any false rejection for any combination of true and false hypotheses.
However, the number of tests can become unfeasible for large values of $m$. A dramatic shortcut is given by selecting the maximum of the test statistics as combining function and is referred to as the max-$T$ method. There are roughly two versions of the max-$T$ method by Westfall and Young (Westfall and Young, 1993; Westfall and Troendle, 2008; Meinshausen et al., 2011): the single-step method and the sequential method. The single-step method is simpler and faster, while the sequential method is more powerful. The single-step approach, based on the matrix $\bm{M}^{n}$ of test statistics, is defined as follows. Here we formulate a version that employs two-sided tests. For every $1\leq j\leq w$, let $m_{j}$ be the maximum of the test statistics $|\bm{M}^{n}_{j,l}|$, $1\leq l\leq m$. Let $m_{(1)},\dots,m_{(w)}$ be the sorted values $m_{1},\dots,m_{w}$. Then the multiple testing method rejects all hypotheses with index $l$ for which $|\bm{M}^{n}|_{1,l}>m_{(\lceil(1-\alpha)w\rceil)}$. The sequential max-$T$ method is defined as follows; after the first step defined above is completed, the procedure is continued in a step-down way. We remove from the matrix $\bm{M}^{n}$ all rows corresponding to the hypotheses rejected in the first step, then we reapply the same procedure described above. This process is continued until there are no more rejections. Note that the procedure with standardized test statistic is defined in the same way.
The following theorem states that the single-step and sequential max-$T$ methods provide strong asymptotic FWER control. Write ${\rm FWER}_{n}$ to indicate potential dependence of the FWER on $n$.
@@
For both the single-step and sequential max-$T$ method, $\limsup_{n\rightarrow\infty}({\rm FWER}_{n})\leq\alpha.$
@@
For both the single-step and the sequential max-$T$ method, the argument is as follows. Recall that $\bm{M}^{n}$ converges in distribution to $\bm{M}$. Let $\bm{M}_{\mathcal{N}}$ be the submatrix of $\bm{M}$ that only contains the rows corresponding to the true null hypotheses. If the matrix $\bm{M}$ is used as input for the multiple testing procedure, then the rows of the submatrix $\bm{M}_{\mathcal{N}}$ are exchangeable, i.e. swapping rows does not change the distribution of the matrix. This implies that the maxT method provides strong FWER control (see e.g. Goeman and Solari, 2010, Theorem 2).
If instead $\bm{M}^{n}$ is used as input, then FWER control follows from the continuous mapping theorem (Van der Vaart, 1998), since $\bm{M}^{n}$ converges to $\bm{M}$ in distribution. ∎
@@ 4.1 Univariate
Figure 1: Logit model, univariate. On abscissa the Bonferroni-adjusted alpha for the control of FWER at level 5%, i.e. .05/m. The flipscores is always inside che confidence bands, while the likelihood ratio test shows an anti-conservative behavior and the other two tests are conservative
We first show a simulation study for univariate tests in Figure 1. The true model is a logistic regression model with one target covariate $X$ and one nuisance covariate $Z$, whose sample size is $n=50$. We test the target parameter $H_{0}:\beta=0$ setting the true regression parameter $\beta=0.$ Four settings are considered: the top-left has true $\gamma$ (the nuisance parameter) equal to 0 and correlation between $X$ and $Z$ equal to 0, the top-right has true $\gamma$ equal to 1 and correlation between $X$ and $Z$ equal to 0, the bottom-left has true $\gamma$ equal to 0 and correlation between $X$ and $Z$ equal to 0.5 and the bottom-right has true $\gamma$ equal to 1 and correlation between $X$ and $Z$ equal to 0.5. A total of 100 000 simulations have been run. We compare the standardized flipscores approach with three standard parametric competitors, namely the Wald, Score and Likelihood Ratio (LRT) tests. The flipscores test is based on 2 000 random flips (we refer to Hemerik and Goeman (2018) for a discussion about the use of a limited number of random flips). We plot 95% simulation confidence bands for the actual control of the significance level. The $x$ axis represents the significance level as a function of the number of hypotheses tested, while the $y$ axis represents the ratio between the Empirical type I error and the nominal level $\alpha$. A Bonferroni correction is applied to the number of hypotheses tested. Note that this plot aims to show the control of the type I error of each testing approach given the actual significance level of the overall analysis due to the number of tests involved.
We can observe that the Flipscores is generally inside the 95% simulation confidence bands. The Wald test is very conservative, the parametric score test is slightly conservative while the likelihood ratio test does not control false positives. More simulations can be find in the appendix for the different settings with a sample size of $n=100$, where we observe similar results.
@@ 4.2 Multivariate
Figure 2: FWER for the multivariate logit model. The flipscores approach control the FWER in each setting. Instead the Wald and score tests are conservative whilst the likelihood ratio test is inflated at low correlations and conservative at high correlations.
Figure 3: Power for the multivariate logit model. The Flipscores approach is more powerful with respect to the other approaches as the correlation increases.
Figures 2 and 3 represent a multivariate simulation. We set a total of 1 000 dependent variables. For 20% they have $\beta=1$ and remaining 80 % have $\beta=0$. It means that we 1 000 logistic regression model, one for each response, which are modelled as function of one common target covariate $X$ with parameter $\beta$ and one common nuisance covariate $Z$ with parameter $\gamma$. Four settings are considered: the top-left has true $\gamma$ equal to 0 and correlation between $X$ and $Z$ equal to 0, the top-right has true $\gamma$ equal to $-1$ and correlation between $X$ and $Z$ equal to 0, the bottom-left has true $\gamma$ equal to 0 and correlation between $X$ and $Z$ equal to 0.5 and the bottom-right has true $\gamma$ equal to $-1$ and correlation between $X$ and $Z$ equal to 0.5.
For controlling the FWER we apply the proposed method with the standardized flipscores, using 2 000 random flips, while the other parametric competitors are corrected with the Bonferroni-Holm procedure.
The $x$ axis of both figures represents the average observed correlation between the dependent variables. Figure 2 represents the empirical FWER of the true null hypothesis. The flipscores is closed to the nominal level, the LRT is anticonservative for low correlation, while the other two methods are highly conservative. Figure 3 illustrates the empirical power for the true alternatives. The flipscores approach has the greatest power, especially for higher levels of correlation, as the other methods do not take advantage of the correlation structure. More simulations can be find in the appendix for the different settings with a sample size of $n=100$, where we observe similar results.
@@ 5 Conclusion
In this paper we focus on the control of the FWER and on the power properties of the multiple testing procedure. However, we can do further inference; indeed, the procedure based on the sign-flip test can be easily extended in order to estimate a lower bound for the true discovery proportion (TDP), that is, the proportion of non-null coefficients over a pre-specified group of hypotheses, using a permutation-based framework (Goeman and Solari, 2011; Blain et al., 2022; Andreella et al., 2023; Vesely et al., 2023).